Required information

Exercise 17-10 Efficiency and profitability analysis LO P3

[The following information applies to the questions displayed below.]

Simon Company’s year-end balance sheets follow.

At December 31 2017 2016 2015
Assets
Cash $ 31,800 $ 35,500 $ 38,200
Accounts receivable, net 89,900 63,500 50,500
Merchandise inventory 112,000 81,400 52,500
Prepaid expenses 10,800 9,250 4,000
Plant assets, net

278,000

252,500 226,000
Total assets $ 522,500 $ 442,150 $ 371,200
Liabilities and Equity
Accounts payable $ 129,600 $ 74,750 $ 51,000
Long-term notes payable secured by
mortgages on plant assets
97,000 100,000 81,200
Common stock, $10 par value 164,000 164,000 164,000
Retained earnings 131,900 103,400 75,000
Total liabilities and equity $ 522,500 $ 442,150 $ 371,200


The company’s income statements for the years ended December 31, 2017 and 2016, follow.

For Year Ended December 31 2017 2016
Sales $ 705,000 $ 530,000
Cost of goods sold $ 415,950 $ 328,600
Other operating expenses 197,400 132,500
Interest expense 12,100 12,800
Income taxes 9,400 8,750
Total costs and expenses 634,850 482,650
Net income $ 70,150 $ 47,350
Earnings per share $ 4.28 $ 2.89


Evaluate the company's efficiency and profitability by computing the following for 2017 and 2016.

(1) Profit margin ratio.

(2) Total asset turnover.

(3) Return on total assets.

Profit Margin Ratio
Choose Numerator: / Choose Denominator: = Profit Margin Ratio
/ = Profit margin ratio
2017 / = %
2016 / = %
Total Asset Turnover
Choose Numerator: / Choose Denominator: = Total Asset Turnover
/ = Total asset turnover
2017 / = times
2016 / = times
Return On Total Assets
Choose Numerator: / Choose Denominator: = Return On Total Assets
/ = Return on total assets
2017 / = %
2016 / = %
Answer #1

Profit margin ratio

Choose numerator / Choose denominator = Profit margin ratio
Net income / Sales = Profit margin ratio
2017 70150 / 705000 = 9.95%
2016 47350 / 530000 = 8.93%

Asset turnover

Choose numerator / Choose denominator = Assets turnover
Sales / Average assets = Assets turnover
2017 705000 / 482325 = 1.46 Times
2016 530000 / 406675 = 1.30 Time

Return on assets

Choose numerator / Choose denominator = Return on assets
Net income / Average assets = Return on assets
2017 70150 / 482325 = 14.54%
2016 47350 / 406675 = 11.64%